OpenStudy (anonymous):

Solve for x: 2x2 - 4x - 14 = 0....(completing the square) I have the answer up to 2(x-1)^2√7 plzz tell me if I'm right and if I'm not plz explan.

OpenStudy (callisto):

Probably, it's not right... 2x^2 - 4x - 14 = 0 2(x^2 - 2x - 7) = 0 x^2 - 2x - 7 = 0 x^2 -2x +(1-1) -7 = 0 (x^2 -2x+1) -1 -7 = 0 (x-1)^2 -8 = 0 (x-1)^2 = 8 Can you continue?

OpenStudy (anonymous):

yes, I think so, thank u very much :D

OpenStudy (callisto):

You're welcome :)

OpenStudy (anonymous):

so for the last part I put them both under √'s?.

OpenStudy (callisto):

That is the second last part.

OpenStudy (anonymous):

okay, I'm kinda stuck..what is the first to the last part.

OpenStudy (callisto):

Take square root on both sides.

OpenStudy (anonymous):

I already said that and u said that waz the 2nd to the last part?

OpenStudy (callisto):

What do you get for that step?

OpenStudy (anonymous):

I get x-1 = + or _ √8

OpenStudy (anonymous):

than you solve for both sides right?

OpenStudy (callisto):

Yes, you need to isolate x on the left..

OpenStudy (anonymous):

for my answer I get x=-8 and x=8?

OpenStudy (callisto):

No!

OpenStudy (anonymous):

x=7 and x=-9?

OpenStudy (callisto):

No! \[x-1 = \pm \sqrt8\] Add 1 to both sides

OpenStudy (anonymous):

soo -1 +8?

OpenStudy (anonymous):

omg , I know this is so simply but I'm just not understanding it srry.

OpenStudy (callisto):

No..... you cannot remove the square root. Moreover, you should add 1, instead of minus 1

OpenStudy (anonymous):

ohh , so x=1!!

OpenStudy (callisto):

NO!!!!!!

OpenStudy (anonymous):

just tell me the answer and then explan it to mee! ....

OpenStudy (callisto):

\[x-1=\pm \sqrt{8}\]Add 1 to both sides\[(x-1)+1=(\pm \sqrt{8})+1\]\[x=1 \pm \sqrt{8}\]

OpenStudy (anonymous):

soo 1-2√2?

OpenStudy (anonymous):

no it's x - 1 = ±2√2? say somthing....!

OpenStudy (anonymous):

heyy person viewing this do you know the answer?

OpenStudy (mimi_x3):

\[2x^{2} -4x-14=0\] \[2\left(\left(x-\frac{2}{2}\right)^{2}-\left(\frac{2}{2}\right)^{2}-7\right) =0\] \[2\left(\left(x-1\right)^{2}-(1)-7\right) = 0 \] \[2\left(\left(x-1\right)^{2}-8\right) =0 \] \[2\left(x-1\right)^{2} - 16 = 0 \] \[2\left(x-1\right)^{2} =16 \] \[\left(x-1\right)^{2} = 8 => \left(x-1\right) = \pm2\sqrt{2} => => x = 1\pm2\sqrt{2} \]

OpenStudy (anonymous):

ty :)